OpenStudy (anonymous):

If the equation of a quadratic is y = 6x2 -4x - 1, then the equation of the axis of symmetry is: x= -1/3 x= 1/3 x=3 x=-3

OpenStudy (mathstudent55):

\(y = ax^2 + bx + c\) \(y' = 2ax + b\) \(2ax + b = 0\) \(2ax = -b\) \(x = -\dfrac{b}{2a} \) Now use the a and b of your problem to find x.

OpenStudy (anonymous):

so its -1/3?

OpenStudy (anonymous):

\[y=6\left( x ^{2}-\frac{ 4 }{6 }x \right)-1\] \[y=6\left( x ^{2} -\frac{ 1 }{ 3 }x+\frac{ 1 }{ 9 }-\frac{ 1 }{9 }\right)-1\] \[y=6\left( x-\frac{ 1 }{ 3 } \right)^{2}-\frac{ 6 }{ 9 }-1\] \[y=6\left( x-\frac{ 1 }{ 3 } \right)^{2}-\frac{ 5 }{ 3 }\] \[axis is x-\frac{ 1 }{3 }=0,or x=\frac{ 1 }{3 }\]

OpenStudy (mathstudent55):

\(x = -\dfrac{b}{2a} \) \(x = -\dfrac{(-4)}{2(6)} \) \(x = \dfrac{4}{12} \) \( x = \dfrac{1}{3} \)

OpenStudy (anonymous):

ohh ok that makes sense now

OpenStudy (mathstudent55):

great